A random sample of 45 cars in the drive-thru of a popular fast food restaurant revealed an average bill of $19.39 per car. The population standard deviation is $5.56. Estimate the mean bill for all cars from the drive-thru with 91% confidence. Round intermediate and final answers to two decimal places.
Answer #1
sample mean 'x̄= 19.390
sample size    n= 45.00
std deviation σ= 5.560
std error ='σx=σ/√n= 0.8288
for 91 % CI value of z= 1.70
margin of error E=z*std error = 1.41
lower bound=sample mean-E= 17.98
Upper bound=sample mean+E= 20.80
from above 91% confidence interval for population mean =(17.98,20.80)